Sometimes probability statements are expressed in terms of odds. The odds in favor of an event A is the following ratio. P (A) P (A) POnot A) PCA For instance, if P(A) 0.60, then P(AC) 0.40 and the odds in favor of A are 0.60 written as 3 to 2 or 3:2. 0.40 4 2 (a) Show that if we are given the odds in favor of event A as n:m, the probability of event A is given by the following P (A) n m Hint: Solve the following equation for PCA). Do this on paper. Your instructor may ask you to turn in this work.) P(A) m POA) (b) A telemarketing supervisor tells a new worker that the odds of making a sale on a single call are 6 to 17. What is the probability of a successful call? (Enter your answer to 2 decimal places.) 3.83 (c) A sports announcer says that the odds a basketball player will make a free throw shot are 4 to 5. What is the probability the player will make the shot? (Enter your answer to 2 decimal places 5.25
Answer #1

(b) probability of succss = 6/(6+17) = 6/23 =0.26

(c)probability of the shot = 4/(4+5) = 4/9 = 0.44